You have invested $250 on the stock market. And have a value rather of $280 after 6 months. What simple interest did your investment cash?( express your answer in percentage)
Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

there are 5 consecutive odd nos. difference between the square of average of 1st two and square of average to last two digit is 564 , find the smallest odd no.
-----------
Use n-4, n-2, n, n+2, n+4 as the numbers.
---
(n-3)^2 = (n+3)^2 - 564
n^2 - 6x + 9 = n^2 + 6n + 9 - 564
-12n = -564
n = 47
---------
n-4 = 43

Answer by fractalier(6550)

Your interest was $30 on a principal of $250.
The formula for simple interest is
I = Prt
Here we look for rate, r. So for a time of half a year,
r = I/(Pt) = $30/(250*.5) = 24% interest